OpenStudy (anonymous):

If g(x) = 2x4 + 3x3 - x2, find g(-1)

OpenStudy (anonymous):

just fill in -1 for x ---> 2*(-1)^4+3*(-1)^3-(-1)^2 ---> 2*1+3*(-1)-1=-2

OpenStudy (anonymous):

g(\(\color{red}{ \sf x }\))=2\(\color{red}{ \sf x }\)^4 + 3\(\color{red}{ \sf x }\)^3 - \(\color{red}{ \sf x }\)^2 g\(\color{red}{ \sf (-1) }\)=2\(\color{red}{ \sf (-1) }\)^4 + 3\(\color{red}{ \sf (-1) }\)^3 - \(\color{red}{ \sf (-1) }\)^2 = 2 - 3 - 1 = -2